Find the area of the rectangles whose sides are:
2 m and 70 cm

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

l = 2 m and b = 70 cm = 0.70 m
Area = l × b = 2 × 0.70
= 1.40 m^2

Thanks